OpenStudy (anonymous):

Find constants a and b so that the minimum for the parabola f(x)=x^2+ax+b is at the point (6,7)

OpenStudy (anonymous):

I thought the derivative equals 2x+a

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I did it right in the last equation.

OpenStudy (anonymous):

\[f'(x)= 2 x +a\\ f(6) = 6^2 + 6 a + b =7\\ f'(6)= 2(6) + a =0\\ \] You find a =12 and b=43

OpenStudy (anonymous):

ok ill try it thank u

OpenStudy (anonymous):

it says b was 43 but a wasnt twelve. Do you know what could have been wrong?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

a=-12 b=43

OpenStudy (anonymous):

ughh thanks hah ah i posted a new one very similar cuz it made me start with new values